OpenStudy (anonymous):

1. Consider the following 3 systems of equations. Which 2 of them are equivalent with one-another (a) 2x - 2y + z = 9 x + y + z = 2 y + z = 1

OpenStudy (anonymous):

figure out what x, y and z are then all will be obvious

OpenStudy (anonymous):

Is there some kind of way to figure them out instead of just using trial and error?

OpenStudy (anonymous):

substitution and addition are two good methods for instance, figure out what y is equal to x + y + z = 2 y = 2 - x - z now substitute y in for a diff. equation: y + z = 1 or 2 - x - z +z = 1 2 - x = 1 x = 1 now you have the value of x. depending on the pattern of your expressions, it could be easier to substitute a value in, or add the two together.

OpenStudy (anonymous):

Oh I get it I think thanks a lot.

OpenStudy (anonymous):

You would need to solve for x,y, and z. After solving for those you can plug it back into the three equations to see which two come out with the same results. If the results are the same then the two equations are equal.